# Numerical investigation of buoyancy-aided mixed convective flow past a square cylinder inclined at 45 degrees

### v3

@@ Abstract
The present study numerically investigates the two-dimensional mixed convective flow of air past a square cylinder placed at an angle of incidence, $\alpha=45^{\circ}$ to the free-stream. We perform direct numerical simulations (DNS) for a Reynolds number (Re) of 100, a range of Richardson numbers (Ri) between 0.0 and 1.0, and a Prandtl number (Pr) of 0.7. The critical Richardson number at which the near-field becomes a steady flow from an unsteady one lies between $0.65$ and $0.7$, along with a simultaneous emergence of the far-field unsteadiness. There is no range of Ri for which the entire flow field is seen to be steady. At a relatively moderate Ri, the flow field reveals the presence of vorticity inversion through the momentum addition in the downstream region. We discuss the dual wake-plume nature of the flow beyond the cylinder. The wake exhibits characteristics similar to those of a buoyant jet in the far-field at increased buoyancy. We explore the cause of the far-field unsteadiness and discuss the mechanism of the observed flow physics using instantaneous and time-averaged flow fields. The important flow quantities, such as force coefficients, vortex shedding frequency, and Nusselt number, are discussed at various Richardson numbers.
@@ I Introduction
It is now well-known that bluff body flows reveal a series of transitions. At Re $\approx$ 5, the flow past a square cylinder at an angle of incidence $\alpha=0^{\circ}$ separates at the front sharp edges, and the two separating shear layers merge downstream of the cylinder forming a symmetric steady recirculation region, whose length increases with an increase in Re. The flow undergoes a Hopf bifurcation (steady to unsteady flow) at Re $\approx$ 45 [1, 2]. Depending on Re, these two shear layers sometimes reattach on the side walls before finally separating at the rear edges of the cylinder. For Re in the range (150-175), a second (spatial) transition takes the two-dimensional flow towards three-dimensionality. Further, an increase in Re takes the flow to turbulence along the Ruelle-Takens-Newhouse route to chaos, as shown by Robichaux et al. [3, 4]. All computations in this study are two-dimensional, as the transition to three-dimensionality occurs well beyond Re $=150$, while our study is for Re $=100$. Remarkably, despite geometrical differences, the wake characteristics of square and circular cylinders exhibit striking similarities [5, 6, 7, 8].
In the case of $\alpha=45^{\circ}$, the Hopf bifurcation occurs at Re $=42$ due to the lower bluffness compared to the $\alpha=0^{\circ}$ case [9]. This is the lowest critical Reynolds number among all angles of incidence in the range $0^{\circ}<\alpha<45^{\circ}$. The second transition occurs around Re $\approx$ 130 as reported by Yoon et al. [10] where there is an onset of three-dimensionality in the flow. Once again, this critical value is the lowest among all angles of incidence in the range $0^{\circ}<\alpha<45^{\circ}$. While the sequence of flow transitions remains identical between $0^{\circ}$ and $45^{\circ}$, the critical Re for each transition is significantly lower in the $\alpha=45^{\circ}$ case.
@@
The heating of the obstacle in such flows leads to numerous new phenomena. In mixed convective flows, the heated cylinder generates density differences in the fluid, adding momentum to the near-field that would otherwise have a momentum deficit due to the presence of the cylinder, which ”suppresses” vortex shedding [18, 19, 20]. Sharma and Eswaran reported that the flow field is steady beyond a critical Richardson number (Ri) of 0.15 [21, 18, 22]. They have also studied the effect of buoyancy-aided (Ri $>0.0$) and buoyancy-opposed (Ri $<0.0$) configurations, where the cylinder is at a higher or lower temperature than the free-stream, respectively. Studies have also investigated flow in the regimes of forced and natural convection [23, 24] as well as natural convection flows with other geometries, which reveal a similar route to chaos as that of forced flows past cylinders [25, 26, 27]. The effect of angle of incidence has been studied for isothermal [9, 10] and non-isothermal flows [28, 29]. There is an effect of the downstream length of the domain on the flow physics, as shown by Dushe [18], who reported that the increased strength of the buoyancy far downstream causes previously unexamined physical phenomena to take place. Most studies have not considered a domain longer than $50d$ ($d$ is the projected length of the cylinder), leading to the loss of far-field flow physics. Motivated by the work of Dushe, we have considered an exceptionally long domain of $110d$ downstream of the cylinder. For ease of further discussion, we have divided the domain beyond the cylinder into three distinct regions: (i) near-field (Y $<15$), (ii) intermediate-field ($15\leq$ Y $\leq 60$), and (iii) far-field (Y $>60$).
@@
Most studies on mixed convective flows past bluff bodies have dealt with the near-field [30, 28, 29, 22]. There has been no numerical investigation on the far-field characteristics of mixed convective flows past cylinders except that of Dushe [18], who has performed a detailed far-field investigation for the $\alpha=0^{\circ}$ case and found further transitions in the far-field at higher Ri. Only one study exists for the $\alpha=45^{\circ}$ case in the buoyancy-aided regime, by Arif and Hasan, [28] who have discussed only the near-field flow. An experimental study by Kimura and Bejan [31] provides qualitative evidence of unsteadiness in the far-field of a natural convection plume resulting from a point heat source. The heat is increased progressively, and changes in the plume structure in the far-field are qualitatively examined. For $\alpha=45^{\circ}$, no work exists that discusses the physics of both the near-field and far-field flow in the buoyancy-aided flow regime and, thus, is the motivation to perform this study to observe the far-field characteristics as well as examine the effect of the cylinder orientation on the nature of the downstream flow.
@@
To achieve the objectives outlined above, we have conducted direct numerical simulations (DNS) of two-dimensional, unconfined, buoyancy-aided (vertically upward flow with gravity acting downward) mixed convective flow of air past a heated square cylinder (heated to a constant wall temperature higher than the free-stream temperature) at a fixed incidence angle, $\alpha=45^{\circ}$ to the free-stream. The present investigation considers flow in the buoyancy-aided regime as the buoyancy force assists the flow in the upward direction. The flow parameters used are Re $=100$ and $0.0\leq Ri\leq 1.0$. Air has been considered as the fluid, the Prandtl number (Pr) of which is taken as 0.7.
@@ II.2 Governing Differential Equations and Boundary Conditions
[…]
@@
$Gr$ is the Grashof number, $g$ is the acceleration due to gravity, $\beta$ is the coefficient of volumetric thermal expansion, $Pr$ is the Prandtl number of the fluid, $\nu$ is the kinematic viscosity of the fluid, $\rho$ is the density of the fluid, $\mu$ is the dynamic viscosity of the fluid, $C_{ph}$ is the specific heat of the fluid, and $K$ is the thermal conductivity of the fluid. $\delta_{i,j}$ is the Kronecker delta.
@@
[…]
At the outlet ($Y=L$), selecting an appropriate outflow boundary condition is challenging because there is no unique prescription for the outflow boundary of external flows. It has been shown that the selection of the outflow boundary condition greatly affects the upstream flow, particularly in subsonic flow. Therefore, it should be chosen to have minimal effect on upstream flow. Accordingly, a convective boundary condition, first prescribed by Orlanski [32], has been employed for both velocity and temperature:
[…]
Where, $v_{c}$ is the average convective velocity of vortices leaving the domain. An optimum value of $v_{c}$ = 0.8 has been used for all calculations performed in the present study.
@@
(a) Complete domain ($-25\leq$ X $\leq 25$, $-10\leq$ Y $\leq 110$)
(b) Near the cylinder vicinity ($-0.85\leq$ X $\leq 0.85$, $-1\leq$ Y $\leq 1$)
@@ II.3 Solution Methodology
[…]
@@
[…]
@@
[…]
@@
Table 2: Comparison of mean Nusselt number and drag coefficient at Re $=100$ ($\alpha=0^{\circ}$).
Richardson number (Ri)
@@
Present Sohankar et al. [9] Yoon et al. [10] Present Sohankar et al. [9]
50 1.63 1.55 1.59 0.144 0.140
100 1.77 1.71 1.72 0.184 0.178
150 1.86 1.85 1.79 0.190 0.194
Table 4: Comparison of mean Nusselt number at Re $=100$ and Ri $=0.0$ ($\alpha=45^{\circ}$).
Reynolds number (Re)
@@
The accuracy of the computed solution depends on the discretization error, which is controlled by the grid resolution. To ensure that the grid is sufficiently fine to reduce the discretization error, a grid independence test has been carried out using three different grid sizes, namely, G1 (619 $\times$ 320), G2 (777 $\times$ 396), and G3 (883 $\times$ 504) at Re $=100$ and Ri $=1.0$ by comparing the same parameters used in the validation. The parameters computed using the three grid sizes, along with the minimum grid spacing in the vicinity of the cylinder, are presented in Table 5. The domain size is kept constant for the grid independence test (120 $\times$ 50). We observe that G2 produces results with minimal deviation from G3. The mean drag coefficient reveals an error of $2.82\%$ and $1.23\%$ for G1 and G2, respectively, with respect to G3. The corresponding error in the case of the Strouhal number is $2.09\%$ and $5.88\%$, and for the mean Nusselt number, it is $1.09\%$ and $0.55\%$, respectively. Given the moderate computational cost and closeness of results with G3, the grid G2 is utilized for all subsequent calculations.
Table 5: Mean drag coefficient, mean Nusselt number and Strouhal number for different grid resolutions at Re $=100$ and Ri $=1.0$.
@@
As mentioned in Section I, integral parameters depend on the blockage ratio. To ensure that the calculated integral parameters are free from dependence on the blockage ratio, a domain independence test has been carried out using two different domains, namely D1 ($120\times 50$, $\beta=2\%$) and D2 ($120\times 67$, $\beta=1.5\%$) at Re $=100$ and Ri $=1.0$ by comparing the same parameters that are used in the validation. Both domains have the same streamwise extent. As the original domain is already very long, it is expected that an increase in its length downstream of the cylinder will not change the flow physics. A detailed investigation on the effect of the domain length has been conducted by Dushe [18]. The mean drag coefficient ($\overline{C}_{D}$), mean Nusselt number ($\overline{Nu}$), and Strouhal number (St) for the domain D1 are 2.44, 5.47, and 0.221. The corresponding values for the domain D2 are 2.45, 5.44, and 0.237. The mean drag coefficient reveals an error of $0.4\%$ for D1 with respect to D2, whereas for mean Nusselt number and Strouhal number, it is $0.6\%$, and $6.8\%$, respectively. The domain independence test reveals that the results do not depend on blockage. Therefore, domain D1 is used for all subsequent calculations.
@@ III.1 Instantaneous Flow
Instantaneous vorticity contours for Re $=100$ and $0.0\leq$ Ri $\leq 1.0$ in the full domain, near-field, and far-field are presented in Figures 3, 4 and 5 respectively. Instantaneous temperature contours in the full domain and near-field are shown in Figures 6 and 7 respectively. All instantaneous contours are phase-synchronized (phase in the shedding cycle when the lift coefficient is minimum) to ensure uniformity in flow visualization.
@@
The instantaneous vorticity contours (see Figure 3) reveal the presence of the Bénard-von Kármán vortex street, signifying the shedding of vortices in the near-field for Ri $=0.0$ (forced convection). There is negative vorticity in the left and positive vorticity in the right half, which spans the entire domain downstream. The two oppositely-signed shear layers attached to the cylinder interact, leading to the shedding of vortices. The flow becomes steady at the far-field due to viscous diffusion. When Ri is increased to a non-zero positive value, it is seen that the sign of vorticity starts as that of a forced flow but slowly inverts beyond a critical distance downstream and persists with this inverted sign for the rest of the domain beyond the critical distance. Additionally, the far-field flow is no longer steady, which was the case for Ri $=0.0$. At higher Ri (0.4-0.7), the inversion of vorticity takes place closer to the cylinder, and the width between the two shear layers increases downstream. The shape of the shed vortices also differs greatly with increasing Ri, which requires further investigation. For Ri $\geq$ 0.7, the near-field flow becomes steady with the suppression of vortex shedding due to the effect of buoyancy, and a large-scale undulation of the shear layers is seen in the far-field region. The shear layer becomes discontinuous, forming isolated patches of vorticity due to the instability developing within the stronger shear layer. As Ri increases, the inception point of the far-field unsteadiness moves closer to the cylinder. The instantaneous vorticity in the cylinder vicinity (see Figure 4) shows regular vortex shedding at Ri $=0.0$ with vortices of the opposite sign being shed alternatively into the wake. With increasing Ri, the transverse extent of the interaction of the positively-signed shear layer with the oppositely-signed shear layer reduces gradually, appearing as though their mutual interaction is being inhibited. For Ri $>$ 0.7, there is minimal interaction between the oppositely-signed shear layers, and they form a separation bubble, similar to forced flows of a sub-critical Re at which the transition to unsteadiness occurs. Vortex shedding from the cylinder does not happen for Ri $>$ 0.7, and the point of inversion of vorticity comes so close to the cylinder that it can be visualized in the near-field plots of vorticity as well (see Figures 4(e) and 4(f)).
(a) Ri $=0.0$
(b) Ri $=0.2$
(c) Ri $=0.4$
(d) Ri $=0.6$
(e) Ri $=0.7$
(f) Ri $=1.0$
@@
Figure 3: Instantaneous vorticity contours in the complete domain for Re $=100$ and $0.0\leq$ Ri $\leq 1.0$: (a) Ri $=0.0$, (b) Ri $=0.2$, (c) Ri $=0.4$, (d) Ri $=0.6$, (e) Ri $=0.7$, (f) Ri $=1.0$.
(a) Ri $=0.0$
(b) Ri $=0.2$
(c) Ri $=0.4$
(d) Ri $=0.6$
(e) Ri $=0.7$
(f) Ri $=1.0$
@@
Figure 4: Instantaneous vorticity contours in the near-field for Re $=100$ and $0.0\leq$ Ri $\leq 1.0$: (a) Ri $=0.0$, (b) Ri $=0.2$, (c) Ri $=0.4$, (d) Ri $=0.6$, (e) Ri $=0.7$, (f) Ri $=1.0$; ($\omega_{baro,min},\omega_{baro,max},\Delta\omega_{baro})\equiv(-0.1,0.1,0.01)$. The flood represents total vorticity, while the lines represent baroclinic vorticity.
(a) Ri $=0.0$
(b) Ri $=0.2$
(c) Ri $=0.4$
(d) Ri $=0.6$
(e) Ri $=0.7$
(f) Ri $=1.0$
@@
Figure 5: Instantaneous vorticity contours in the far-field for Re $=100$ and $0.0\leq$ Ri $\leq 1.0$: (a) Ri $=0.0$, (b) Ri $=0.2$, (c) Ri $=0.4$, (d) Ri $=0.6$, (e) Ri $=0.7$, (f) Ri $=1.0$; ($\omega_{baro,min},\omega_{baro,max},\Delta\omega_{baro})\equiv(-0.05,0.05,0.005)$. The flood represents total vorticity while the lines represent baroclinic vorticity.

### v4

@@ Abstract
The present study numerically investigates the two-dimensional mixed convective flow of air past a square cylinder placed at an angle of incidence $\alpha=45^{\circ}$ to the free-stream. We perform direct numerical simulations (DNS) for a Reynolds number (Re) of 100, a range of Richardson numbers (Ri) between 0.0 and 1.0, and a Prandtl number (Pr) of 0.7. The critical Richardson number at which the near-field becomes a steady flow from an unsteady one lies between $0.65$ and $0.7$, along with a simultaneous emergence of the far-field unsteadiness. There is no range of Ri for which the entire flow field is seen to be steady. At a relatively moderate Ri, the flow field reveals the presence of vorticity inversion through the momentum addition in the downstream region. We discuss the dual wake-plume nature of the flow behind the cylinder. The wake exhibits characteristics similar to those of a plume, revealing a self-similar behavior in the far-field at increased buoyancy. We explore the cause of the far-field unsteadiness and discuss the mechanism of the observed flow physics using instantaneous and time-averaged flow fields. The important flow quantities, such as force coefficients, vortex shedding frequency, and Nusselt number, are discussed at various Richardson numbers.
preprint: The following article has been accepted for publication in Physical Review Fluids.
@@ I Introduction
It is now well-known that bluff body flows reveal a series of transitions. At Re $\approx$ 5, the flow past a square cylinder at an angle of incidence $\alpha=0^{\circ}$ separates at the front sharp edges, and the two separating shear layers merge downstream of the cylinder forming a symmetric steady recirculation region, whose length increases with an increase in Re. The flow undergoes a Hopf bifurcation (steady to unsteady flow) at Re $\approx$ 45 [1, 2]. Depending on Re, these two shear layers sometimes reattach on the side walls before finally separating at the rear edges of the cylinder. For Re in the range (150-175), a second (spatial) transition takes the two-dimensional flow towards three-dimensionality. Further, an increase in Re takes the flow to turbulence along the Ruelle-Takens-Newhouse route to chaos, as shown by Robichaux et al. [3, 4]. Remarkably, despite geometrical differences, the wake characteristics of square and circular cylinders exhibit striking similarities [5, 6, 7, 8].
In the case of $\alpha=45^{\circ}$, the Hopf bifurcation occurs at Re =42=42 due to the lower bluffness compared to the $\alpha=0^{\circ}$ case [9]. This is the lowest critical Reynolds number among all angles of incidence in the range $0^{\circ}<\alpha<45^{\circ}$. Upto an Re of 100, there exist five distinct regimes of the two-dimensional flow with the formation of the main wake along with sub-wakes around the trailing downstream corner of the cylinder [10]. Due to the symmetric shape of the obstacle about the centerline, the wake possesses spatio-temporal symmetry, wherein the flow reflects about the centerline after each half shedding period [11], similar to the $\alpha=0^{\circ}$ case. The second transition occurs around Re $\approx$ 120 at which the flow becomes three-dimensional, as reported by Jiang [11] and Yoon et al. [12]. Once again, this critical value is the lowest among all angles of incidence in the range $0^{\circ}<\alpha<45^{\circ}$. While the sequence of flow transitions remains identical between $0^{\circ}$ and $45^{\circ}$, the critical Re for each transition is significantly lower in the $\alpha=45^{\circ}$ case. All computations in this study are two-dimensional, as the transition to three-dimensionality occurs around Re =120=120, while our study is for Re =100=100.
@@
The heating of the obstacle in such flows leads to numerous new phenomena. In mixed convective flows, the heated cylinder generates density differences in the fluid, adding momentum to the near-field that would otherwise have a momentum deficit due to the presence of the cylinder, which ”suppresses” vortex shedding [20, 21, 22]. Sharma and Eswaran reported that the flow field is steady beyond a critical Richardson number (Ri) of 0.15 [23, 20, 24]. They have also studied the effect of buoyancy-aided (Ri >0.0>0.0) and buoyancy-opposed (Ri <0.0<0.0) configurations, where the cylinder is at a higher or lower temperature than the free-stream, respectively. Studies have also investigated flow in the regimes of forced and natural convection [25, 26] as well as natural convection flows with other geometries, which reveal a similar route to chaos as that of forced flows past cylinders [27, 28, 29]. The effect of angle of incidence has been studied for isothermal [9, 12] and non-isothermal flows [30, 31, 32, 33]. The effect of heating level, free-stream inclination [34], and Prandtl number [35] has been studied for square cylinders, and similar studies exist for elliptical cylinders [36]. Many works have concentrated on the mixed convective flow in the cross-buoyancy configuration with emphasis on the effect of baroclinic vorticity on the wake structures [37, 38, 39]. There is an effect of the downstream length of the domain on the flow physics, as shown by Dushe [20], who reported that the increased strength of the buoyancy far downstream causes previously unexamined physical phenomena to take place. Most studies have not considered a domain longer than $50d$ ($d$ is the projected length of the cylinder), leading to the loss of far-field flow physics. Motivated by the work of Dushe, we have considered an exceptionally long domain of $110d$ downstream of the cylinder. For ease of further discussion, we have divided the domain beyond the cylinder into three distinct regions: (i) near-field (Y <15<15), (ii) intermediate-field ($15\leq$ Y $\leq 60$), and (iii) far-field (Y >60>60).
@@
Natural convective plumes occur in a variety of industrial and environmental situations [40]. Studying the plume behaviour, which occurs far downstream of the source, is key in order to characterize such phenomena. Most studies on mixed convective flows past bluff bodies have dealt with the near-field [41, 30, 31, 24, 32, 33, 34]. There has been no numerical investigation on the far-field characteristics of mixed convective flows past cylinders except that of Dushe [20], who has performed a detailed far-field investigation for the $\alpha=0^{\circ}$ case and found further transitions in the far-field at higher Ri. Very few studies exist for the $\alpha=45^{\circ}$ case in the buoyancy-aided regime [30, 32, 33], where only the near-field flow has been discussed. An experimental study by Kimura and Bejan [42] provides qualitative evidence of unsteadiness in the far-field of a natural convection plume resulting from a point heat source. The heat is increased progressively, and changes in the plume structure in the far-field are qualitatively examined. For $\alpha=45^{\circ}$, no work exists that discusses the physics of both the near-field and far-field flow in the buoyancy-aided flow regime and, thus, is the motivation to perform this study to observe the far-field characteristics as well as examine the effect of the cylinder orientation on the nature of the downstream flow.
@@
To achieve the objectives outlined above, we have conducted direct numerical simulations (DNS) of two-dimensional, unconfined, buoyancy-aided (vertically upward flow with gravity acting downward) mixed convective flow of air past a heated square cylinder (heated to a constant wall temperature higher than the free-stream temperature) at a fixed incidence angle, $\alpha=45^{\circ}$ to the free-stream. The present investigation considers flow in the buoyancy-aided regime as the buoyancy force assists the flow in the upward direction. The flow parameters used are Re =100=100 and $0.0\leq Ri\leq 1.0$. Air has been considered as the fluid, the Prandtl number (Pr) of which is taken as 0.7.
@@ II.2 Governing Differential Equations and Boundary Conditions
$\frac{\partial u_{i}}{\partial x_{i}}=0$ (1)
$\frac{\partial u_{i}}{\partial\tau}+\frac{\partial(u_{j}u_{i})}{\partial x_{j}}=-\frac{\partial p}{\partial x_{i}}+\frac{1}{Re}\left[\frac{\partial^{2}u_{i}}{\partial x_{i}^{2}}\right]+Ri\theta\delta_{i2}$ (2)
$\frac{\partial\theta}{\partial\tau}+\frac{\partial(u_{i}\theta)}{\partial x_{i}}=\frac{1}{RePr}\left[\frac{\partial^{2}\theta}{\partial x_{i}^{2}}\right]$ (3)
@@
GrGr is the Grashof number, $g$ is the acceleration due to gravity, $\beta$ is the coefficient of volumetric thermal expansion, $Pr$ is the Prandtl number of the fluid, $\nu$ is the kinematic viscosity of the fluid, $\rho$ is the density of the fluid, $\mu$ is the dynamic viscosity of the fluid, $C_{ph}$ is the specific heat of the fluid, and $K$ is the thermal conductivity of the fluid. $\delta_{i,j}$ is the Kronecker delta.
@@
$\partialderivative{v}{x}=0,\hskip 5.0ptu=0$
$\partialderivative{\theta}{x}=0$
At the outlet ($Y=L$), selecting an appropriate outflow boundary condition is challenging because there is no unique prescription for the outflow boundary of external flows. It has been shown that the selection of the outflow boundary condition greatly affects the upstream flow, particularly in subsonic flow. Therefore, it should be chosen to have minimal effect on upstream flow. It has been shown in the past that the convective boundary condition (CBC) is superior to other boundary conditions such as zero Neumann boundary condition (NBC) or vanishing second derivatives (VSD) by Sohankar et al. [9, 43] and others [44]. In particular, Sohankar et al. [9] have shown that the CBC reduces computational time, iterations for convergence, and the time taken to reach the periodic flow. Moreover, the CBC allows for the smooth passage of vortices out of the domain, while the NBC and VSD conditions distort and smear vortices exiting the domain. Accordingly, the convective boundary condition, first prescribed by Orlanski [45], has been employed for both velocity and temperature:
$\partialderivative{u_{i}}{t}+v_{c}\partialderivative{u_{i}}{y}=0$
$\partialderivative{\theta}{t}+v_{c}\partialderivative{\theta}{y}=0$
Where, $v_{c}$ is the average convective velocity of vortices leaving the domain. An optimum value of $v_{c}$ = 0.8, which provides a faster convergence without any loss of accuracy, has been used for all calculations performed in the present study.
@@
(a) Complete domain ($25\leq$ X $\leq 25$, $10\leq$ Y $\leq 110$)
(b) Near the cylinder vicinity ($0.85\leq$ X $\leq 0.85$, $1\leq$ Y $\leq 1$)
@@ II.3 Solution Methodology
$\frac{u_{i}^{*}-u_{i}^{n}}{\Delta t}=\left[\frac{3}{2}G(u_{i},u_{j})^{n}-\frac{1}{2}G(u_{i},u_{j})^{n-1}\right]-\partialderivative{p^{n}}{x_{i}}$ (6)
@@
$p^{\prime}=\dfrac{-\omega_{0}\partialderivative{u_{i}^{*}}{x_{i}}}{2\Delta t\left[\dfrac{1}{(\Delta x)^{2}}+\dfrac{1}{(\Delta y)^{2}}\right]}$ (7)
@@
$\implies Nu=\dfrac{hd}{K}=\left.-\partialderivative{\theta}{n}\right|_{n=0}$
@@
Table 2: Comparison of mean Nusselt number and drag coefficient at Re =100=100 ($\alpha=0^{\circ}$).
Richardson
number (Ri)
@@
Present
Sohankar
et al. [9]
Yoon
et al. [12]
Yadav
et al. [10]
Present
Sohankar
et al. [9]
Yadav
et al. [10]
50 1.63 1.55 1.59 1.64 0.144 0.140 0.141
100 1.77 1.71 1.72 1.78 0.180 0.178 0.180
150 1.86 1.85 1.79 1.94 0.190 0.194 0.200
Table 4: Comparison of mean Nusselt number at Re =100=100 and Ri =0.0=0.0 ($\alpha=45^{\circ}$).
Reynolds
number (Re)
@@
The accuracy of the computed solution depends on the discretization error, which is controlled by the grid resolution. To ensure that the grid is sufficiently fine to reduce the discretization error, a grid independence test has been carried out using three different grid sizes, namely, G1 (619 $\times$ 320), G2 (777 $\times$ 396), and G3 (883 $\times$ 504) at Re =100=100 and Ri =1.0=1.0 by comparing the same parameters used in the validation. The parameters computed using the three grid sizes, along with the minimum grid spacing in the vicinity of the cylinder, are presented in Table 5. The domain size is kept constant for the grid independence test (120 $\times$ 50). We observe that G2 produces results with minimal deviation from G3. The mean drag coefficient reveals an error of $2.82\%$ and $1.23\%$ for G1 and G2, respectively, with respect to G3. The corresponding error in the case of the Strouhal number is $2.09\%$ and $5.88\%$, and for the mean Nusselt number, it is $1.09\%$ and $0.55\%$, respectively. Given the moderate computational cost and closeness of results with G3, the grid G2 is utilized for all subsequent calculations.
Table 5: Mean drag coefficient, mean Nusselt number and Strouhal number for different grid resolutions at Re =100=100 and Ri =1.0=1.0.
@@
As mentioned in Section I, integral parameters depend on the blockage ratio. To ensure that the calculated integral parameters are free from dependence on the blockage ratio, a domain independence test has been carried out using two different domains, namely D1 ($120\times 50$, $\beta=2\%$) and D2 ($120\times 67$, $\beta=1.5\%$) at Re =100=100 and Ri =1.0=1.0 by comparing the same parameters that are used in the validation. Both domains have the same streamwise extent. As the original domain is already very long, it is expected that an increase in its length downstream of the cylinder will not change the flow physics. A detailed investigation on the effect of the domain length has been conducted by Dushe [20]. The mean drag coefficient ($\overline{C}_{D}$), mean Nusselt number ($\overline{Nu}$), and Strouhal number (St) for the domain D1 are 2.44, 5.47, and 0.221. The corresponding values for the domain D2 are 2.45, 5.44, and 0.237. The mean drag coefficient reveals an error of $0.4\%$ for D1 with respect to D2, whereas for the mean Nusselt number and Strouhal number, it is $0.6\%$, and $6.8\%$, respectively. The domain independence test reveals that the results do not depend on blockage. Therefore, domain D1 is used for all subsequent calculations.
In order to prevent numerical instabilities caused by the use of an explicit scheme for temporal discretization in the present study, the time increment is determined by subjecting it to two restrictions, namely, the Courant–Friedrichs–Lewy (CFL) condition, and the grid Fourier number (Fo) restriction corresponding to convection and diffusion dominated regions respectively. We have calculated the minimum time step throughout the domain as prescribed by the above two criteria. Further, the lower value between the two conditions is multiplied by a factor less than unity (between 0.2 and 0.4) to ensure stability for the second-order accurate spatially and temporally discretized equations used in the present study. The typical value of the time step is found to be 0.0005, which was chosen after performing a time-step independence study.
@@ III.1 Instantaneous Flow
Instantaneous vorticity contours for Re =100=100 and $0.0\leq$ Ri $\leq 1.0$ in the full domain, near-field, and far-field are presented in Figures 3, 4 and 5 respectively. Instantaneous temperature contours in the full domain and near-field are shown in Figures 6 and 7 respectively. All instantaneous contours are phase-synchronized (phase in the shedding cycle when the lift coefficient is minimum) to ensure uniformity in flow visualization.
@@
The instantaneous vorticity contours (see Figure 3) reveal the presence of the Bénard-von Kármán vortex street, signifying the shedding of vortices in the near-field for Ri =0.0=0.0 (forced convection). There is negative vorticity in the left and positive vorticity in the right half, which spans the entire domain downstream. The two oppositely signed shear layers attached to the cylinder interact, leading to the shedding of vortices. The flow becomes steady at the far-field due to viscous diffusion. When Ri is increased to a non-zero positive value, it is seen that the sign of vorticity starts as that of a forced flow but slowly inverts beyond a critical distance downstream and persists with this inverted sign for the rest of the domain beyond the critical distance. Additionally, the far-field flow is no longer steady, which was the case for Ri =0.0=$0.$ At higher Ri (0.4-0.7), the inversion of vorticity takes place closer to the cylinder, and the width between the two shear layers increases downstream. The shape of the shed vortices also differs greatly with increasing Ri, which requires further investigation. For Ri $\geq$ 0.7, the near-field flow becomes steady with the suppression of vortex shedding due to the effect of buoyancy, and a large-scale undulation of the shear layers is seen in the far-field region. The shear layer becomes discontinuous, forming isolated patches of vorticity due to the instability developing within the stronger shear layer. As Ri increases, the inception point of the far-field unsteadiness moves closer to the cylinder. The instantaneous vorticity in the cylinder vicinity (see Figure 4) shows regular vortex shedding at Ri =0.0=0.0 with vortices of the opposite sign being shed alternatively into the wake. With increasing Ri, the transverse extent of the interaction of the positively-signed shear layer with the oppositely-signed shear layer reduces gradually, appearing as though their mutual interaction is being inhibited. For Ri >> 0.7, there is minimal interaction between the oppositely-signed shear layers, and they form a separation bubble, similar to forced flows of a sub-critical Re at which the transition to unsteadiness occurs. Vortex shedding from the cylinder does not happen for Ri >> 0.7, and the point of inversion of vorticity comes so close to the cylinder that it can be visualized in the near-field plots of vorticity as well (see Figures 4(e) and 4(f)).
(a) Ri =0.0=0.0
(b) Ri =0.2=0.2
(c) Ri =0.4=0.4
(d) Ri =0.6=0.6
(e) Ri =0.7=0.7
(f) Ri =1.0=1.0
@@
Figure 3: Instantaneous vorticity contours in the complete domain for Re =100=100 and $0.0\leq$ Ri $\leq 1.0$: (a) Ri =0.0=0.0, (b) Ri =0.2=0.2, (c) Ri =0.4=0.4, (d) Ri =0.6=0.6, (e) Ri =0.7=0.7, (f) Ri =1.0=1.0; ($\omega_{min},\omega_{max},\Delta\omega)\equiv(-0.5,0.5,0.01)$.
(a) Ri =0.0=0.0
(b) Ri =0.2=0.2
(c) Ri =0.4=0.4
(d) Ri =0.6=0.6
(e) Ri =0.7=0.7
(f) Ri =1.0=1.0
@@
Figure 4: Instantaneous vorticity contours in the near-field for Re =100=100 and $0.0\leq$ Ri $\leq 1.0$: (a) Ri =0.0=0.0, (b) Ri =0.2=0.2, (c) Ri =0.4=0.4, (d) Ri =0.6=0.6, (e) Ri =0.7=0.7, (f) Ri =1.0=1.0; ($\omega_{min},\omega_{max},\Delta\omega)\equiv(-0.5,0.5,0.01)$, ($\omega_{baro,min},\omega_{baro,max},\Delta\omega_{baro})\equiv(-0.1,0.1,0.01)$. The flood represents total vorticity, while the lines represent baroclinic vorticity.
(a) Ri =0.0=0.0
(b) Ri =0.2=0.2
(c) Ri =0.4=0.4
(d) Ri =0.6=0.6
(e) Ri =0.7=0.7
(f) Ri =1.0=1.0
@@
Figure 5: Instantaneous vorticity contours in the far-field for Re =100=100 and $0.0\leq$ Ri $\leq 1.0$: (a) Ri =0.0=0.0, (b) Ri =0.2=0.2, (c) Ri =0.4=0.4, (d) Ri =0.6=0.6, (e) Ri =0.7=0.7, (f) Ri =1.0=1.0; ($\omega_{min},\omega_{max},\Delta\omega)\equiv(-0.5,0.5,0.01)$, ($\omega_{baro,min},\omega_{baro,max},\Delta\omega_{baro})\equiv(-0.05,0.05,0.005)$. The flood represents total vorticity while the lines represent baroclinic vorticity.